# ContextGPT: Infusing LLMs Knowledge into Neuro-Symbolic Activity Recognition Models

### v1

@@ 6. Conclusions and Future Work
We have several plans for future work. First, in this work we considered a general LLM incorporating knowledge from many domains. We want to investigate how to specialize the LLM in the HAR domain. This is challenging since it requires identifying reliable sources to effectively fine-tune the model.
Besides fine-tuning, we will investigate personalization aspects. Different individuals may have customized habits requiring ad-hoc models (e.g., “Bob usually enjoys running at the park some late afternoons after work, even if it’s raining”). Thus, we will investigate how to introduce into the prompt personalized aspects that may help in capturing personalized relationships between activities and contexts.
Finally, ContextGPT currently generates a list of activities consistent with the current context without considering fuzziness. However, associating a consistency score with each activity may lead to more effective knowledge infusion (Arrotta et al., 2022). The main challenge in this direction is that LLMs are generally not good at dealing with numbers and mathematical operations (Frieder et al., 2023).
Finally, we will also investigate alternative LLM models (e.g., GPT-4, LLAMA2) and NeSy approaches (e.g., the Semantic Loss method recently proposed in (Arrotta et al., 2023a)) and their impact on the recognition rate.
Acknowledgments
The authors want to thank Mohcen Laaroussi for his excellent work on software implementation. This work was partially supported by Italian Ministry of University and Research with project “MUSA-Multilayered Urban Sustainability Action” (project ID: ECS_00000037).

### v2

@@ VI Conclusions and Future Work
We have several plans for future work. First, in this work we considered a general LLM incorporating knowledge from many domains. We want to investigate how to specialize the LLM in the HAR domain. This may be obtained with RAG or fine-tuning techniques. Our investigation also shed some light on the context information released to external LLM services that may expose the subjects to privacy threats. We plan to investigate if more lightweight and open-source LLMs running on trusted edge machines may provide the same knowledge infusion quality.
Acknowledgements
The authors thank Mochen Laaroussi for his excellent implementation work. This work was supported in part by projects FAIR (PE00000013) and SERICS (PE00000014) under the NRRP MUR program funded by the EU - NGEU. Views and opinions expressed are however those of the authors only and do not necessarily reflect those of the European Union or the Italian MUR. Neither the European Union nor the Italian MUR can be held responsible for them.